use elimination to solve the system 3x+5y=16 and 8x-5y=28 for x
Answer by tutorcecilia(2144)

3x+5y=16
8x-5y=28 [perform the indicated operation]
____________
11x+0y=44
.
11x=44 [solve for x]
x=44/11=4 [plug-in either equation and solve for y]
.
3(4)+5y=16
12+5y=16
5y=16-12
y=4/5
.
check by plugging the values of x and y into either equation
3x+5y=16
3(4)+5(4/5)=16
12+4=16
16=16